# NDKF: A Neural-Enhanced Distributed Kalman Filter for Nonlinear Multi-Sensor Estimation

### v1

@@ Abstract
We propose a Neural-Enhanced Distributed Kalman Filter (NDKF) for multi-sensor state estimation in nonlinear systems. Unlike traditional Kalman filters that rely on explicit, linear models and centralized data fusion, the NDKF leverages neural networks to learn both the system dynamics and measurement functions directly from data. Each sensor node performs local prediction and update steps using these learned models and exchanges only compact summary information with its neighbors via a consensus-based fusion process, which reduces communication overhead and eliminates a single point of failure. Our theoretical convergence analysis establishes sufficient conditions under which the local linearizations of the neural models guarantee overall filter stability and provides a solid foundation for the proposed approach. Simulation studies on a 2D system with four partially observing nodes demonstrate that the NDKF significantly outperforms a distributed Extended Kalman Filter. These outcomes, as yielded by the proposed NDKF method, highlight its potential to improve the scalability, robustness, and accuracy of distributed state estimation in complex nonlinear environments.
@@ I Introduction
Distributed Kalman filtering has long been explored to enable multi-sensor state estimation without relying on a centralized fusion node. Early work on consensus-based and federated filters [1, 2, 3] allowed individual nodes to update local estimates and share compact summaries–such as means and covariances–with their neighbors. While effective for linear or mildly nonlinear systems, these methods require careful coordination to avoid overconfidence and inconsistent estimates [4]. To handle nonlinearities and model uncertainties, extensions of the standard Kalman filter have been developed. Variants of the Extended Kalman Filter (EKF) [5] approximate local Jacobians to manage nonlinear dynamics, while unscented Kalman filters (UKF) [6] employ deterministic sampling to capture higher-order moments without linearization. Both approaches have been adapted for distributed settings [7, 8], though they depend on explicit system models that may not fully capture the complexities of real-world dynamics. Robust techniques such as covariance intersection methods [9, 10] further address issues like outliers and unknown correlations, yet they still hinge on predefined models.
In parallel, the integration of machine learning into filtering frameworks has led to neural-based Kalman filters [11, 12]. Rather than relying solely on hand-crafted models, these approaches train neural networks to learn state transitions or measurement mappings directly from data. Notably, KalmanNet [13] incorporates a recurrent neural network into the Kalman filter to learn the Kalman gain, which results in overcoming limitations associated with mismatches in noise statistics and model inaccuracies. However, most neural filtering approaches have been applied in centralized settings or under the assumption of linear models at individual nodes. Moreover, incorporating neural predictions into a distributed filtering framework introduces several challenges: ensuring stability and convergence despite potential biases or large local errors, managing partial or heterogeneous measurements across multiple distributed nodes, and balancing the computational cost of repeated neural network evaluations and Jacobian computations against real-time requirements.
@@
    A hybrid filtering framework that integrates neural network approximations of both system dynamics and measurement functions into the distributed Kalman filtering process to address complex nonlinear behaviors.
@@
    Integration of a consensus-based distributed fusion mechanism into the filtering process to enable scalable, decentralized state estimation, accompanied by a detailed computational complexity analysis.
@@
    A rigorous stability and convergence analysis for the NDKF under local linearization and Lipschitz conditions.
@@
    Validation of the proposed approach on a 2D multi-sensor scenario, demonstrating the efficacy of the NDKF compared to a conventional EKF baseline.
@@ II-A System Setup
$\mathbf{x}_{k+1}=f_{\theta}\bigl{(}\mathbf{x}_{k}\bigr{)}+\mathbf{w}_{k},$ (1)
where $f_{\theta}(\cdot)$ is a (potentially nonlinear) function approximated by a neural network with parameters $\theta$, and $\mathbf{w}_{k}$ is process noise, assumed to be zero-mean with covariance matrix $\mathbf{Q}\succeq\mathbf{0}$. We employ a network of $N$ sensor nodes, each collecting noisy measurements of the state. Node $i\in\{1,2,\dots,N\}$ obtains:
$\mathbf{y}_{k,i}=h_{\theta,i}\bigl{(}\mathbf{x}_{k}\bigr{)}+\mathbf{v}_{k,i},$ (2)
where $h_{\theta,i}(\cdot)$ is the local measurement function for node $i$, potentially also represented or assisted by a neural network, and $\mathbf{v}_{k,i}$ is measurement noise with covariance $\mathbf{R}_{i}\succeq\mathbf{0}$. Depending on the application, each $h_{\theta,i}(\cdot)$ may observe a subset of the state or a particular nonlinear transformation of $\mathbf{x}_{k}$.
@@ II-B Neural Network for System and Measurement Modeling
where $(\mathbf{x}_{k}^{(j)},\mathbf{x}_{k+1}^{(j)})$ pairs are drawn from a training dataset of $M$ samples. The minimization of (3) ensures the network accurately models the underlying system transitions.
@@ II-C Kalman Filter with Neural Network Integration
The Kalman filter is traditionally defined for linear dynamics and observation models. In this work, we preserve the filter’s two-stage prediction and update structure but explicitly replace the system and measurement models with neural-network-based functions $f_{\theta}$ and $h_{\theta,i}$. At each time step $k$, the filter uses these learned models for state propagation and local measurement updates, respectively.
@@ II-C1 Prediction Step
Let $\hat{\mathbf{x}}_{k|k}$ and $\mathbf{P}_{k|k}$ denote the estimated mean and covariance of the state at time $k$ (after processing the measurements up to $k$). The prediction step for time $k{+}1$ uses the learned dynamics function $f_{\theta}$ from (1):
@@ II-C2 Local Measurement Update
$\mathbf{S}_{k,i}\;=\;\mathbf{H}_{\theta,i,k}\,\mathbf{P}_{k|k-1}\,\mathbf{H}_{\theta,i,k}^{T}\;+\;\mathbf{R}_{i},$ (7)
$\mathbf{K}_{k,i}\;=\;\mathbf{P}_{k|k-1}\,\mathbf{H}_{\theta,i,k}^{T}\,\mathbf{S}_{k,i}^{-1},$ (8)
$\hat{\mathbf{x}}_{k|k,i}\;=\;\hat{\mathbf{x}}_{k|k-1}\;+\;\mathbf{K}_{k,i}\,\Bigl{[}\mathbf{y}_{k,i}\;-\;h_{\theta,i}\bigl{(}\hat{\mathbf{x}}_{k|k-1}\bigr{)}\Bigr{]},$ (9)
$\mathbf{P}_{k|k,i}\;=\;\bigl{(}\mathbf{I}-\mathbf{K}_{k,i}\,\mathbf{H}_{\theta,i,k}\bigr{)}\,\mathbf{P}_{k|k-1}.$ (10)
In (7)–(10), the subscript $(k|k-1)$ indicates the predicted quantities from (5)–(6), and $\mathbf{R}_{i}$ is the measurement noise covariance at node $i$. If $h_{\theta,i}(\cdot)$ is linear or partially known, $\mathbf{H}_{\theta,i,k}$ can be simplified accordingly.
@@ II-D Distributed Fusion Mechanism
II-E Initialization and Parameter Tuning
@@
II-F Scalability and Computational Complexity
@@
Each fusion round (as in Section II-D) involves inverting or combining $N$ local covariance matrices and summing information vectors. If each node $i$ handles contributions from its neighbors in $\mathcal{N}_{i}$, the total cost scales with $|\mathcal{N}_{i}|$. In the worst-case scenario of a fully connected network, each node processes contributions from $N-1$ other nodes. The matrix operations remain $O(n^{3})$ per node (for $n$-dimensional state), plus the cost of exchanging local estimates.

### v2

@@ Abstract
We propose a Neural-Enhanced Distributed Kalman Filter (NDKF) for multi-sensor state estimation in nonlinear systems. Unlike traditional Kalman filters that rely on explicit analytical models and assume centralized fusion, NDKF leverages neural networks to replace analytical process and measurement models with learned mappings while each node performs local prediction and update steps and exchanges only compact posterior summaries with its neighbors. This distributed design reduces communication overhead and avoids a central fusion bottleneck. We provide sufficient mean-square stability conditions under bounded Jacobians and well-conditioned innovations, together with practically checkable proxies such as Jacobian norm control and innovation monitoring. We also discuss consistency under learned-model mismatch, including covariance inflation and covariance-intersection fusion when cross-correlations are uncertain. Simulations on a 2D nonlinear system with four partially observing nodes show that NDKF outperforms a distributed EKF baseline under model mismatch and yields improved estimation accuracy with modest communication requirements.
@@ I Introduction
Distributed Kalman filtering has long been explored to enable multi-sensor state estimation without relying on a centralized fusion node. Early work on consensus-based filters [1] and [2], and on federated approaches [3], allowed individual nodes to update local estimates and share compact summaries (such as means and covariances) with their neighbors. While effective for linear or mildly nonlinear systems, these methods require careful coordination to avoid overconfidence and inconsistent estimates [4].
To handle nonlinearities and model uncertainties, extensions of the standard Kalman filter have been developed. Variants of the Extended Kalman Filter (EKF) [5] approximate local Jacobians to manage nonlinear dynamics, with established stochastic stability conditions under boundedness assumptions [6]. Unscented Kalman filters (UKF) [7] employ deterministic sampling to capture higher-order moments without linearization. Both approaches have been adapted for distributed settings [8] and [9], though they depend on explicit system models that may not fully capture the complexities of real-world dynamics. Robust fusion techniques such as covariance intersection [10, 11] address unknown cross-correlations and guard against overconfident fusion; they are complementary to, rather than substitutes for, local outlier rejection mechanisms.
In parallel, the integration of machine learning into filtering frameworks has led to neural-based Kalman filters. Examples include deep latent state-space models [12] and [13], and KalmanNet-style gain learning [14]. The broader lineage also includes dual Kalman filtering ideas [15] and recent multiagent nonlinear filtering-and-learning analyses [16]. However, most neural filtering approaches are applied in centralized settings or assume linear models at individual nodes. Extending them to distributed, partially observed multi-agent systems raises concrete challenges: ensuring stability despite learned-model bias, preventing overconfidence under unknown cross-correlations, and managing the cost of repeated neural evaluations and Jacobians in real time.
NDKF targets these gaps. (i) Each node learns its own measurement mapping while the process mapping can be learned from shared logs or simulation; nodes exchange only posterior summaries. (ii) Information-form consensus provides a principled fusion route and admits a robust covariance-intersection fallback when cross-correlations are unknown. (iii) Our analysis supplies sufficient stability conditions under local linearization with Lipschitz-bounded neural Jacobians and introduces practically checkable proxies (e.g., innovation conditioning and Jacobian norm control) that can be verified or enforced during deployment. We do not assume parameter sharing; here $\theta$ is a collective placeholder for all trainable parameters (process and per-node measurement), and only posterior summaries (not $\theta$) are communicated.
@@
    An integrated filtering framework combining neural network approximations of system dynamics and measurement functions with distributed Kalman filtering to effectively address complex nonlinear behaviors.
@@
    A consensus-based distributed fusion mechanism to enable scalable, decentralized state estimation, complemented by a detailed computational complexity analysis and an optional covariance-intersection fallback when cross-correlations are unknown.
@@
    A mean-square stability and convergence analysis for the NDKF under local linearization and Lipschitz continuity, together with practically checkable proxies for deployment.
@@
    Validation through simulation studies on a 2D multi-sensor scenario, demonstrating superior performance of the proposed NDKF compared to a conventional distributed EKF baseline.
Practical Training Without Ground-Truth States
A common concern is that learning $f_{\theta}$ and $h_{\theta,i}$ requires state labels that are unavailable at deployment. In NDKF, we use regimes that avoid this requirement:
  • Supervised via proxies: brief calibration runs with external instrumentation or high-fidelity simulators (“digital twins”) provide approximate state labels to pretrain $f_{\theta}$ and $h_{\theta,i}$.
  • Latent/self-supervised: treat $\mathbf{x}_{k}$ as latent and train by maximizing predictive likelihood or minimizing multi-sensor innovation residuals over sequences, leveraging [12], [13], and [14] as building blocks. No state labels are required.
  • Residual learning: start from a nominal model $f_{0},h_{0,i}$ and learn small corrections $\Delta f,\Delta h_{i}$ using pseudo-labels from high-trust smoothers or consensus posteriors; regularize the residuals to preserve stability.
These regimes keep training feasible when states are hidden, and align with the distributed nature of our method: nodes train $h_{\theta,i}$ locally, while $f_{\theta}$ can be learned centrally or from simulated traces and then deployed.
Representative Deployment Scenarios
NDKF is particularly appealing in settings where sensors are spatially distributed, measurements are heterogeneous, and accurate first-principles models are unavailable. Examples include cooperative ground or aerial robots with complementary sensing, environmental monitoring with mixed acoustic/range/chemical sensors, intelligent transportation systems with roadside and onboard measurements, and industrial process monitoring across multiple stations. In such deployments, the main practical challenges are communication delays or packet drops, time synchronization, sensor bias drift, occasional outliers or faults, and limited on-node compute. The present design addresses the communication burden by exchanging posterior summaries rather than raw data, and it can be combined with innovation gating and covariance inflation to improve resilience under field conditions [17].
@@ II-A System Setup
$\mathbf{x}_{k+1}=f_{\theta_{f}}\bigl(\mathbf{x}_{k},\mathbf{u}_{k}\bigr)+\mathbf{w}_{k},$ (1)
where $f_{\theta_{f}}(\cdot,\cdot)$ is a potentially nonlinear function approximated by a neural network with parameters $\theta_{f}$, $\mathbf{u}_{k}$ denotes an optional control input, and $\mathbf{w}_{k}$ is zero-mean process noise with covariance matrix $\mathbf{Q}\succeq\mathbf{0}$. In the benchmark considered in Section IV, the system is autonomous and we therefore set $\mathbf{u}_{k}=\mathbf{0}$ for simplicity.
We employ a network of $N$ sensor nodes, each collecting noisy measurements of the state. Node $i\in\{1,2,\dots,N\}$ obtains
$\mathbf{y}_{k,i}=h_{\theta_{h,i},i}\bigl(\mathbf{x}_{k}\bigr)+\mathbf{v}_{k,i},$ (2)
where $h_{\theta_{h,i},i}(\cdot)$ is the local measurement function for node $i$, represented or assisted by a neural network with parameters $\theta_{h,i}$, and $\mathbf{v}_{k,i}$ is zero-mean measurement noise with covariance $\mathbf{R}_{i}\succeq\mathbf{0}$. Depending on the application, each $h_{\theta,i}(\cdot)$ may observe a subset of the state or a particular nonlinear transformation of $\mathbf{x}_{k}$. Note that the parameter sets $\theta$ used in $f_{\theta}(\cdot)$ and $h_{\theta,i}(\cdot)$ are not necessarily the same and can be independently trained neural networks with distinct parameters.
Here, the process noise $\mathbf{w}_{k}$ and the measurement noise $\mathbf{v}_{k,i}$ are assumed to be zero-mean and temporally white (i.e., uncorrelated across time steps). Moreover, these noise processes are mutually independent, with covariances $\mathbf{Q}$ and $\mathbf{R}_{i}$, respectively.
@@
Remark 1
(Parameterization, control inputs, and communication). We use $\theta_{f}$ for the dynamics network and $\theta_{h,i}$ for the measurement network at node $i$; no parameter sharing is assumed unless explicitly stated. When a nominal affine control channel is known, one may also write
$\mathbf{x}_{k+1}=f_{\theta_{f}}(\mathbf{x}_{k})+\mathbf{B}_{k}\mathbf{u}_{k}+\mathbf{w}_{k}.$
During filtering, nodes exchange only posterior summaries (means/covariances or information pairs), not raw measurements and not neural parameters.
@@ II-B Neural Network for System and Measurement Modeling
Here, the training samples $\{(\mathbf{x}_{k}^{(j)},\mathbf{x}_{k+1}^{(j)})\}_{j=1}^{M}$ are obtained from historical measurements or high-fidelity simulations. The minimization of (3) ensures the network accurately models the underlying system transitions. In scenarios where the true system dynamics are unknown, one may employ an approximate model or direct sensor observations during an initial phase to generate these samples and to enable the neural network to learn a corrective (residual) representation of the dynamics.
@@
To mitigate distribution shift, we (i) learn residual corrections on top of nominal models when available, (ii) regularize networks for smoothness (spectral/weight decay) to control Jacobians, (iii) use early stopping and multiple random initializations to avoid poor local minima, and (iv) enable lightweight online adaptation of small residual heads if the operating regime drifts, while keeping the core model fixed to preserve stability.
Prediction use $f_{\theta_{f}}$ and $\mathbf{F}_{\theta_{f},k}$ compute $\hat{\mathbf{x}}_{k+1|k},\ \mathbf{P}_{k+1|k}$ Local update at node $i$ use $h_{\theta_{h,i},i}$ and $\mathbf{H}_{\theta_{h,i},i,k}$ compute $\hat{\mathbf{x}}_{k+1|k+1,i},\ \mathbf{P}_{k+1|k+1,i}$ Information packaging $\mathbf{W}_{i}=\mathbf{P}_{k+1|k+1,i}^{-1}$ $\mathbf{z}_{i}=\mathbf{W}_{i}\hat{\mathbf{x}}_{k+1|k+1,i}$ Neighbor fusion exchange $(\mathbf{W}_{j},\mathbf{z}_{j})$ consensus or covariance intersection Output fused posterior $\hat{\mathbf{x}}_{k+1|k+1,i}^{(\mathrm{fusion})}$ $\mathbf{P}_{k+1|k+1,i}^{(\mathrm{fusion})}$ Input posterior $\hat{\mathbf{x}}_{k|k,i}$ $\mathbf{P}_{k|k,i}$ Shared dynamics NN $f_{\theta_{f}}(\mathbf{x},\mathbf{u})$ 3 hidden layers, 128 units Node-$i$ meas. NN $h_{\theta_{h,i},i}$ Local measurement $\mathbf{y}_{k+1,i}$ Neighbor messages $\{(\mathbf{W}_{j},\mathbf{z}_{j})\}_{j\in\mathcal{N}_{i}}$ Node $i$ runtime pipelineOnly posterior summaries are    exchanged across neighbors  used as input posteriorfor the next step
Figure 1: NDKF overview. Each node executes prediction using the shared dynamics network $f_{\theta_{f}}$, applies a local update with its own measurement network $h_{\theta_{h,i},i}$, packages the posterior into information form $(\mathbf{W}_{i},\mathbf{z}_{i})$, and fuses neighbor messages through consensus or covariance intersection. Raw measurements remain local.
@@ II-C Kalman Filter with Neural Network Integration
The proposed filter preserves the classical two-stage prediction/update but differs in three practical aspects: (i) the state transition and per-node observation functions are learned (and may be nonlinear and heterogeneous across nodes); (ii) Jacobians $\mathbf{F}_{\theta,k}$ and $\mathbf{H}_{\theta,i,k}$ come from autodiff of these networks or finite differences; and (iii) the local posteriors are fused in information form (with an optional CI fallback) to mitigate overconfidence when cross-correlations are unknown.
@@ II-C1 Prediction Step
Let $\hat{\mathbf{x}}_{k|k}$ and $\mathbf{P}_{k|k}$ denote the estimated mean and covariance of the state at time $k$ (after processing the measurements up to $k$). The prediction step for time k+1k{+}1 uses the learned dynamics function $f_{\theta}$ from (1):
@@ II-C2 Local Measurement Update
$\mathbf{S}_{k+1,i}\;=\;\mathbf{H}_{\theta,i,k+1}\,\mathbf{P}_{k+1|k}\,\mathbf{H}_{\theta,i,k+1}^{T}\;+\;\mathbf{R}_{i},$ (7)
$\mathbf{K}_{k+1,i}\;=\;\mathbf{P}_{k+1|k}\,\mathbf{H}_{\theta,i,k+1}^{T}\,\mathbf{S}_{k+1,i}^{-1},$ (8)
$\hat{\mathbf{x}}_{k+1|k+1,i}=\hat{\mathbf{x}}_{k+1|k}+\mathbf{K}_{k+1,i}\,\Bigl[\mathbf{y}_{k+1,i}-h_{\theta,i}\bigl(\hat{\mathbf{x}}_{k+1|k}\bigr)\Bigr],$ (9)
$\mathbf{P}_{k+1|k+1,i}\;=\;\bigl(\mathbf{I}-\mathbf{K}_{k+1,i}\,\mathbf{H}_{\theta,i,k+1}\bigr)\,\mathbf{P}_{k+1|k}.$ (10)
In (7)–(10), the subscript (k+1|k)(k{+}1|k) indicates the predicted quantities from (5)–(6), and $\mathbf{R}_{i}$ is the measurement noise covariance at node $i$. If $h_{\theta,i}(\cdot)$ is linear or partially known, $\mathbf{H}_{\theta,i,k+1}$ can be simplified accordingly.
Relation to distributed EKF
The algebraic prediction/update equations in NDKF intentionally retain the EKF structure; the novelty lies in what is linearized and how local posteriors are fused. Specifically, (i) the process and measurement maps are learned from data rather than assumed analytically known, (ii) each node may carry a distinct learned measurement model, and (iii) fusion is performed in information form with a consistency-preserving covariance-intersection fallback when cross-correlations are uncertain. Consequently, if the learned models are replaced by exact analytical ones and covariance intersection is not needed, NDKF reduces locally to a distributed EKF/information filter. We emphasize that NDKF does not eliminate linearization error; rather, it addresses a different failure mode, namely inaccurate analytical process and measurement models. When linearization error dominates, derivative-free distributed filters may be preferable.
@@ II-D Distributed Fusion Mechanism
When cross-correlations between local posteriors are unknown, (12) can be replaced by covariance intersection (CI) [10, 18]: $\mathbf{W}_{\mathrm{CI}}=\omega\mathbf{W}_{a}+(1-\omega)\mathbf{W}_{b}$, $\mathbf{z}_{\mathrm{CI}}=\omega\mathbf{z}_{a}+(1-\omega)\mathbf{z}_{b}$, $\omega\in[0,1]$. CI guarantees consistency (non-overconfident fusion) at the cost of conservatism and integrates seamlessly with our information-form exchange.
II-E Algorithm 1: One NDKF step at node $i$
Here we summarize the per-time-step procedure executed at node $i$, covering prediction, local measurement update, information-form packaging, and consensus fusion (with an optional covariance-intersection fallback), mapping inputs $\hat{\mathbf{x}}_{k|k},\mathbf{P}_{k|k},\mathbf{y}_{k+1,i}$ and neighbors’ messages to the fused posterior $\hat{\mathbf{x}}_{k+1|k+1,i}^{(\mathrm{fusion})}$ and $\mathbf{P}_{k+1|k+1,i}^{(\mathrm{fusion})}$.
  1. 1.
    Prediction: $\hat{\mathbf{x}}_{k+1|k}\leftarrow f_{\theta}\!\left(\hat{\mathbf{x}}_{k|k}\right)$, $\mathbf{P}_{k+1|k}\leftarrow\mathbf{F}_{\theta,k}\mathbf{P}_{k|k}\mathbf{F}_{\theta,k}^{\!\top}+\mathbf{Q}$, with $\mathbf{F}_{\theta,k}:=\nabla f_{\theta}(\hat{\mathbf{x}}_{k|k})$ (autodiff or finite differences).
  2. 2.
    Local update: form $\mathbf{H}_{\theta,i,k+1}:=\nabla h_{\theta,i}(\hat{\mathbf{x}}_{k+1|k})$, innovation covariance $\mathbf{S}_{k+1,i}$, gain $\mathbf{K}_{k+1,i}$ via (7)–(8), and update (9)–(10).
  3. 3.
    Information packaging: $\mathbf{W}_{i}\leftarrow\mathbf{P}_{k+1|k+1,i}^{-1}$, $\mathbf{z}_{i}\leftarrow\mathbf{W}_{i}\hat{\mathbf{x}}_{k+1|k+1,i}$.
  4. 4.
    Consensus fusion: aggregate $\sum_{j\in\mathcal{N}_{i}}\mathbf{W}_{j},\sum_{j\in\mathcal{N}_{i}}\mathbf{z}_{j}$ to obtain fused $\mathbf{P}_{k+1|k+1,i}^{(\mathrm{fusion})}$ and $\hat{\mathbf{x}}_{k+1|k+1,i}^{(\mathrm{fusion})}$ via (13)–(14). Optional: if cross-correlations are unknown, use covariance intersection in place of (12).
II-F Initialization and Parameter Tuning
@@
II-G Scalability and Computational Complexity
@@
Each fusion round (as in Section II-D) involves inverting or combining $N$ local covariance matrices and summing information vectors. If each node $i$ handles contributions from its neighbors in $\mathcal{N}_{i}$, the total cost scales with $|\mathcal{N}_{i}|$. In the worst-case scenario of a fully connected network, each node processes contributions from N1N-1 other nodes. The matrix operations remain $O(n^{3})$ per node (for $n$-dimensional state), plus the cost of exchanging local estimates.
@@
Per-node compute is dominated by Jacobian evaluations and $O(n^{3})$ factorizations; communication scales with node degree $|\mathcal{N}_{i}|$ (not with $N$ if the graph is sparse). In practice we found that increasing state dimension $n$ calls for moderately wider networks to preserve approximation quality (e.g., hidden width scaling sub-quadratically in $n$), while depth increases yield diminishing returns compared to Jacobian-cost growth. Performance degradation with $N$ is typically sub-linear when degree is bounded; over-smoothing can be avoided by limiting consensus rounds or using CI. These guidelines help select capacity without excessive complexity.